OpenStudy (anonymous):

What is the solution to the equation 49x=sqrt 7^(2x+6)

OpenStudy (jamesj):

49x or 49^x ?

OpenStudy (anonymous):

49^x ,sorry

OpenStudy (jamesj):

Notice first then that because 49 = 7^2, that 49^x = 7^(2x).

OpenStudy (jamesj):

yes? Now square both sides: \[ 7^{4x} = 7^{2x+6} \] making sense?

OpenStudy (anonymous):

sort of

OpenStudy (jamesj):

So first, happy with 49^x = 7^2x ?

OpenStudy (anonymous):

yes i understand that part

OpenStudy (jamesj):

\[ 49^x = (7^2)^x = 7^{2x} \] Now that is equal to \[ \sqrt{7^{2x+6}} = (7^{2x+6})^{1/2} \]

OpenStudy (anonymous):

okay

OpenStudy (jamesj):

and that is equal to \[ 7^{(2x+6)/2} = 7^{x+3} \]

OpenStudy (jamesj):

Hence the original equation \[ 49^x = 7^{2x+6} \] implies \[ 7^{2x} = 7^{x+3} \] because \[ 49^x = 7^{2x} \] and \[ \sqrt{7^{2x+6}} = 7^{x+3} \]

OpenStudy (anonymous):

okay

OpenStudy (jamesj):

the original equation has a square root I dropped.

OpenStudy (jamesj):

So now we have \[ 7^{2x} = 7^{x+3} \] what does that imply?

OpenStudy (anonymous):

umm, idk

OpenStudy (jamesj):

if \[ 7^a = 7^b \] what can you say about a and b?

OpenStudy (anonymous):

they are equal

OpenStudy (jamesj):

Yes, a = b. Hence if \[ 7^{2x} = 7^{x+3} \] what can you say.

OpenStudy (anonymous):

they are equal also

OpenStudy (jamesj):

what's equal? make it explicit; write down the equation.

OpenStudy (jamesj):

\[ 7^a = 7^b \ \implies \ a = b \] hence \[ 7^{2x} = 7^{x+3} \ \implies \ ...what? \]

OpenStudy (jamesj):

still there?

OpenStudy (anonymous):

i need help on this to. so your saying 2x = x +3??

OpenStudy (jamesj):

Yes

OpenStudy (anonymous):

This helped me so much! Thanks for taking the time to answer this for her!